# Improving Genomic Prediction using High-dimensional Secondary Phenotypes: the Genetic Latent Factor Approach

### v1

@@ 2.1. Model and assumptions
These assumptions imply that the genetic covariance matrix $\bm{\Sigma}_{ss}^{g}$ equals the model-implied genetic covariance matrix $\bm{\Sigma}_{ss}^{g}(\bm{\Theta})=\mathbf{\Lambda}\mathbf{\Lambda}^{\top}+\mathbf{\Psi}=\mathbb{E}\left[(\mathbf{\Lambda}\bm{\xi}_{j}+\bm{\varepsilon}_{j})(\mathbf{\Lambda}\bm{\xi}_{j}+\bm{\varepsilon}_{j})^{\top}\right]$ where $\bm{\Theta}=\{m,\mathbf{\Lambda},\mathbf{\Psi}\}$. Therefore, $\bar{\mathbf{y}}_{s(j)}\sim\mathcal{N}_{p}(\mathbf{0},\mathbf{\Lambda}\mathbf{\Lambda}^{\top}+\mathbf{\Psi}+r^{-1}\bm{\Sigma}_{ss}^{\epsilon})$. The covariance of the genetic components $\mathbf{g}_{s(j)}$ can thus be decomposed into common ($\mathbf{\Lambda}\mathbf{\Lambda}^{\top}$) and unique ($\mathbf{\Psi}$) elements. Using standard covariance algebra we can also obtain the joint distribution of the secondary feature BLUEs and latent factors as
@@
(5) $\begin{bmatrix}\bar{\mathbf{y}}_{s(j)}^{\top}\\ \bm{\xi}_{j}\end{bmatrix}\sim\mathcal{N}_{p+m}\left[\mathbf{0},\bm{\Sigma}(\bm{\Theta})_{\xi}^{y}\right]\text{,}$
@@
We will use the model above to project the observed secondary features onto a lower-dimensional latent space, followed by multivariate genomic prediction in seven steps. We first use the plot-level data to obtain estimates for the genetic and residual correlation matrices of the observed features (I). We then redundancy filter the data (II) to ensure there are no pairs of features with near perfect genetic correlations, followed by regularization of the filtered genetic and residual correlation matrices (III). We then estimate the latent dimension (IV) and fit the factor model (V). Next, we obtain estimates for the latent factor scores (VI) and use those to select a subset and calculate multivariate genomic predictions using the standard BLUP equations (VII). The remainder of this section is devoted to the details of these seven steps.
2.2. Covariance matrix estimation (step I)
The factor analytic approach in gfBLUP largely revolves around the genetic and residual covariance matrices of the observed secondary data. We can obtain estimates of these matrices by using REML (restricted maximum likelihood) and the kinship matrix $\mathbf{K}$ for a small number of features. When the number of secondary features $p$ exceeds $5\text{-}10$, however, the computational complexity becomes prohibitive and REML estimates can no longer be obtained [16]. As this is usually the case for high-dimensional HTP data, we use data containing genotypic replicates and correct for any experimental design factors to obtain pseudo-CRD (pseudo completely randomized design) data as a pre-processing step. Note again that in this case, assuming balanced data, the total number of samples is simply the number of genotypes times the number of replicates, i.e., $n=n_{g}r$. We then assume that $\mathbf{Z}\mathbf{K}\mathbf{Z}^{\top}=\mathbf{Z}\mathbf{Z}^{\top}$, i.e., $\mathbf{K}=\mathbf{I}_{n_{g}}$, and obtain estimates for $\bm{\Sigma}_{ss}^{g}$ and $\bm{\Sigma}_{ss}^{\epsilon}$ using sums of squares as measures for between and within-genotype variation [17]:
(6) $\begin{split}\mathbf{MS}_{ss}^{g}&=(\bar{\mathbf{Y}}_{s}-\bar{\mathbf{Y}}_{s}^{g})^{\top}(\bar{\mathbf{Y}}_{s}-\bar{\mathbf{Y}}_{s}^{g})(n_{g}-1)^{-1}\\ \mathbf{MS}_{ss}^{\epsilon}&=(\bar{\mathbf{Y}}_{s}^{g}-\mathbf{Y}_{s})^{\top}(\bar{\mathbf{Y}}_{s}^{g}-\mathbf{Y}_{s})(n-n_{g})^{-1}\\ \hat{\bm{\Sigma}}_{ss}^{g}&=(\mathbf{MS}_{ss}^{g}-\mathbf{MS}_{ss}^{\epsilon})r^{-1}\\ \hat{\bm{\Sigma}}_{ss}^{\epsilon}&=\mathbf{MS}_{ss}^{\epsilon}\text{,}\end{split}$
where $\mathbf{Y}_{s}$ is the $n\times p$ data matrix containing the secondary features. Overall means and genotypic means are located in $\bar{\mathbf{Y}}_{s}$ and $\bar{\mathbf{Y}}_{s}^{g}$, respectively. A well-known issue with the estimation of such covariance matrices using multivariate sums of squares is the fact that sufficient within-genotype variation can result in $\hat{\bm{\Sigma}}_{ss}^{g}$ having negative eigenvalues [18] as it is not a moment matrix but rather the difference of two moment matrices (as opposed to $\hat{\bm{\Sigma}}_{ss}^{\epsilon}$ which is a moment matrix and therefore at worst positive-semidefinite if $p>n$). We deal with this issue by simply finding the nearest positive-definite matrix using Higham’s algorithm [19]. From here on, $\hat{\bm{\Sigma}}_{ss}^{g}$ refers to this nearest positive definite matrix.
We emphasize that the assumption of independent genetic effects is only made for the genetic and residual covariance estimators in equation (6). We use a marker-based kinship matrix in the final genomic prediction step. The assumption of diagonal kinship ($\mathbf{K}=\mathbf{I}_{n_{g}}$) in (6) only means that the subsequent factor analysis is based on the covariance between the total (rather than only the additive) genetic effects. This seems a reasonable approximation since for most quantitative traits the genetic variance is predominantly additive [20].
2.3. Redundancy filtering (step II)
Given the fact that multicollinearity is common in HTP data, some of the $p$ secondary features will likely show near perfect genetic correlation and are therefore redundant. We filter the genetic correlation matrix using the algorithm described by Peeters et al. [21] with an absolute genetic correlation threshold $\tau$ (i.e., after filtering, no two features will have an absolute genetic correlation $|\rho^{g}|$ equal to, or exceeding $\tau$). This filtering step produces $\hat{\mathbf{R}}_{ss}^{g*}\in\mathbb{R}^{p^{*}\times p^{*}}$ containing only the $p^{*}$ secondary features that pass filtering ($p^{*}\leq p$). We subset the residual correlation matrix and data matrix so they too only contain the remaining $p^{*}$ features.
2.4. Regularization (step III)
The filtered estimates $\hat{\mathbf{R}}_{ss}^{g*}$ and $\hat{\mathbf{R}}_{ss}^{\epsilon*}$ may still be ill-conditioned or singular which is why regularization is required. To achieve this, we use penalized maximum likelihood forms in the remainder of the gfBLUP method [21, 22]:
where $\vartheta_{g},\vartheta_{\epsilon}\in(0,1]$. We find the optimal values for the penalty parameters $\vartheta_{g}^{\ddagger}$ and $\vartheta_{\epsilon}^{\ddagger}$ through $K$-fold cross-validation where we split the data by dividing the $n_{g}$ genotypes over $K$ disjoint folds and minimize the following expression using Brent’s algorithm [23]:
where $\left(\mathbf{R}_{ss}^{*}(\vartheta)\right)_{\neg k}$ is the penalized genetic or residual correlation matrix given the corresponding penalty $\vartheta$ estimated using the $K-1$ folds. The non-penalized correlation matrix $(\mathbf{R}_{ss}^{*})_{k}$ is estimated using the left-out fold. Each left-out fold is weighed by $w_{k}$, which is equal to the number of genotypes or individuals in that fold for the genetic and residual correlation matrix, respectively. See Peeters et al. [21] for further details. We use an efficient implementation to minimize (8) which results in iterations of complexity $O(p^{*})$ as opposed to $O(p^{*3})$ due to the repeated inversion of $\mathbf{R}_{ss}^{*}(\vartheta)_{\neg k}$ required in the naive minimization. This leads to faster optimization, especially if optimal penalty values are close to the bounds. Details regarding our implementation can be found in SM Section S.4. Note that both $\hat{\mathbf{R}}_{ss}^{g*}(\vartheta_{g}^{\ddagger})$ and $\hat{\mathbf{R}}_{ss}^{\epsilon*}(\vartheta_{\epsilon}^{\ddagger})$ balance the unbiased but high-variance estimated correlation matrices with the biased but stable identity matrix, optimizing the bias-variance trade-off. We may use another target matrix $\mathbf{T}$ instead of the identity if prior information is available in the form of, for example, covariance estimates from earlier experiments. The only requirement is that $\mathbf{T}$ is a positive-definite correlation matrix.
2.5. Latent dimension (step IV)
In principle, the model specified in (4) is valid for any latent dimension $m\leq\frac{1}{2}\left[2p+1-(8p+1)^{1/2}\right]$ [14]. We must specify an estimate for $m$ prior to fitting the factor model and obtaining an estimate for the other parameters in $\bm{\Theta}$, however. To do so we use a lower eigenvalue bound based on the Marchenko-Pastur distribution [24]:
$\omega_{+}=\left(1+\sqrt{\dfrac{p^{*}}{n_{g}}}\right)^{2}\text{,}$
where $n_{g}$ is the number of genotypes used to estimate the genetic covariance matrix as described in Section 2.2. We then determine $\tilde{m}$:
2.6. Factor analytic modeling (step V)
With regularized estimates for the correlation matrices and latent dimension available, we now proceed with the factor analytic dimension reduction of the redundancy filtered secondary data. In the context of plant breeding, we are mainly interested in genetic correlations and will therefore fit the factor model using a reduced form of the secondary data covariance matrix. Recall that the full covariance of the BLUEs given in Section 2.1 was $\bm{\Sigma}_{ss}^{p*}=\bm{\Sigma}_{ss}^{g*}+r^{-1}\bm{\Sigma}_{ss}^{\epsilon*}$, with stars added to denote the redundancy filtering. The reduced covariance that we use to fit the factor model is thus $\bm{\Sigma}_{ss}^{p*}-r^{-1}\bm{\Sigma}_{ss}^{\epsilon*}=\bm{\Sigma}_{ss}^{g*}=\mathbf{\Lambda}\mathbf{\Lambda}^{\top}+\mathbf{\Psi}$. Given the fact that maximum-likelihood estimation of $\bm{\Theta}$ is scale-free [25], we opt to estimate the loadings and uniquenesses using correlation rather than covariance matrices. Furthermore, we recognize that the sample genetic correlation matrix $\hat{\mathbf{R}}_{ss}^{g*}$ is a sufficient statistic for the population correlation matrix $\mathbf{R}_{ss}^{g*}$, meaning the maximum likelihood estimate $\hat{\bm{\Theta}}$, given a latent dimension $\tilde{m}$, can be found by minimizing the following discrepancy function [26]:
(9) $F\left[\mathbf{R}_{ss}^{g*}(\bm{\Theta});\hat{\mathbf{R}}_{ss}^{g*};\tilde{m}\right]=\ln\Big{|}\mathbf{R}_{ss}^{g*}(\bm{\Theta})\Big{|}+\operatorname{tr}\left[\hat{\mathbf{R}}_{ss}^{g*}\mathbf{R}_{ss}^{g*}(\bm{\Theta})^{-1}\right]-\ln\Big{|}\hat{\mathbf{R}}_{ss}^{g*}\Big{|}-p^{*}\text{,}$
where we constrain $\mathbf{\Lambda}^{\top}\mathbf{\Psi}^{-1}\mathbf{\Lambda}$ to be a diagonal matrix with ordered diagonal elements for identification. We replace $\hat{\mathbf{R}}_{ss}^{g*}$ by the regularized estimate $\hat{\mathbf{R}}_{ss}^{g*}(\vartheta_{g}^{\ddagger})$ as minimizing (9) requires a positive-definite and ideally well-conditioned sample genetic correlation matrix [21]. We employ a Varimax rotation once we have obtained the estimates for $\mathbf{\Lambda}$ and $\mathbf{\Psi}$ to facilitate interpretability [27]. In essence, the varimax rotation attempts to bring the loadings matrix close to having simple structure, i.e., a structure in which variables load highly on a single factor and have near zero loadings on the other factors. To avoid notational clutter we simply use $\hat{\mathbf{\Lambda}}$ to refer to either the rotated or non-rotated estimated loadings. The described factor analytic model is implemented using an efficient matrix-free estimation method [28].
2.7. Estimation of factor scores (step VI)
We would now like to obtain estimates for the factor scores $\bm{\xi}_{j}$. These factor scores are essentially estimates of the locations of genotypes in latent space, based on the many noisy but observable secondary feature measurements. They can thus be viewed as the unobserved phenotypes of a small number of latent traits that generate the observed phenotypes for the many secondary features. We will use the joint distribution of the secondary feature BLUEs and latent factors (5) in a slightly modified Thomson regression approach [29] to obtain $\hat{\bm{\xi}}_{j}$. Recalling that we estimated $\bm{\Theta}$ on the correlation scale, we first convert $\hat{\mathbf{\Lambda}}$ and $\hat{\mathbf{\Psi}}$ back to the covariance scale:
$\hat{\mathbf{\Lambda}}_{\Sigma}=\mathbf{D}\hat{\mathbf{\Lambda}}\text{ and }\hat{\mathbf{\Psi}}_{\Sigma}=\mathbf{D}\hat{\mathbf{\Psi}}\mathbf{D}\text{, where }\mathbf{D}=(\hat{\bm{\Sigma}}_{ss}^{g*}\circ\mathbf{I}_{p^{*}})^{1/2}\text{.}$
@@ 2.7. Estimation of factor scores (step VI)
We can now obtain the matrix $\mathbf{P}$ mapping the redundancy filtered secondary features to their projections on the lower-dimensional latent space through $\mathbb{E}(\bm{\xi}_{j}|\bar{\mathbf{y}}_{s(j)}^{*})$. Given that the joint distribution (5) is normal, we can obtain this conditional expectation as $\bm{\Sigma}(\bm{\Theta})_{\xi y}\left[\bm{\Sigma}(\bm{\Theta})_{yy}\right]^{-1}\bar{\mathbf{y}}_{s(j)}^{*}$ [30]. Using the Woodbury matrix identity [31], we rewrite this last expression to produce:
@@
$\begin{split}\tilde{\bm{\xi}}_{j}&=\Big{[}\mathbf{I}_{\tilde{m}}+\hat{\mathbf{\Lambda}}^{\top}_{\Sigma}(\hat{\mathbf{\Psi}}_{\Sigma}+r^{-1}\hat{\bm{\Sigma}}_{ss}^{\epsilon*})^{-1}\hat{\mathbf{\Lambda}}_{\Sigma}\Big{]}^{-1}\hat{\mathbf{\Lambda}}^{\top}_{\Sigma}(\hat{\mathbf{\Psi}}_{\Sigma}+r^{-1}\hat{\bm{\Sigma}}_{ss}^{\epsilon*})^{-1}\bar{\mathbf{y}}_{s(j)}^{*}\\ &=\mathbf{P}^{\top}\bar{\mathbf{y}}_{s(j)}^{*}\text{.}\end{split}$
@@
(11) $\begin{split}\tilde{\bm{\Xi}}&=\bar{\mathbf{Y}}_{s}^{*}\left[\hat{\mathbf{\Psi}}_{\Sigma}+r^{-1}\hat{\bm{\Sigma}}_{ss}^{\epsilon*}(\vartheta_{\epsilon}^{\ddagger})\right]^{-1}\hat{\mathbf{\Lambda}}_{\Sigma}\left\{\mathbf{I}_{\tilde{m}}+\hat{\mathbf{\Lambda}}_{\Sigma}^{\top}\left[\hat{\mathbf{\Psi}}_{\Sigma}+r^{-1}\hat{\bm{\Sigma}}_{ss}^{\epsilon*}(\vartheta_{\epsilon}^{\ddagger})\right]^{-1}\hat{\mathbf{\Lambda}}_{\Sigma}\right\}^{-1}\\ &=\bar{\mathbf{Y}}_{s}^{*}\mathbf{P}\text{.}\end{split}$
Note that $\text{var}(\bar{\mathbf{y}}_{s(j)}^{*})\rightarrow\bm{\Sigma}_{ss}^{g*}$ as $r\rightarrow\infty$, and as a result the projection $\mathbf{P}$ approaches the matrix that would be obtained using the ordinary Thomson regression approach.

### v3

@@ 2.2.1. Model and assumptions
These assumptions imply that the genetic covariance matrix $\bm{\Sigma}_{ss}^{g}$ equals the model-implied genetic covariance matrix $\bm{\Sigma}_{ss}^{g}(\bm{\Theta})=\mathbf{\Lambda}\mathbf{\Lambda}^{\top}+\mathbf{\Psi}=\mathbb{E}\left[(\mathbf{\Lambda}\bm{\xi}_{j}+\bm{\varepsilon}_{j})(\mathbf{\Lambda}\bm{\xi}_{j}+\bm{\varepsilon}_{j})^{\top}\right]$ where $\bm{\Theta}=\{m,\mathbf{\Lambda},\mathbf{\Psi}\}$. Therefore, $\bar{\mathbf{y}}_{s(j)}\sim\mathcal{N}_{p}(\mathbf{0},\mathbf{\Lambda}\mathbf{\Lambda}^{\top}+\mathbf{\Psi}+r^{-1}\bm{\Sigma}_{ss}^{\epsilon})$. The covariance of the genetic components $\mathbf{g}_{s(j)}$ can thus be decomposed into common ($\mathbf{\Lambda}\mathbf{\Lambda}^{\top}$) and unique ($\mathbf{\Psi}$) elements.
We will use the model above to project the observed secondary features onto a lower-dimensional latent space in six steps.We first use the plot-level data to obtain estimates for the genetic and residual correlation matrices of the observed secondary features (I). We then redundancy filter the secondary data (II) to ensure there are no pairs of features with near perfect genetic correlations, followed by regularization of the filtered genetic and residual correlation matrices (III). We then estimate the latent dimension (IV) and fit the factor model (V). Next, we obtain estimates for the latent factor scores (VI), completing the dimension reduction. The remainder of this section is devoted to the details of these six steps.
2.2.2. Covariance matrix estimation (step I)
The factor analytic approach in glfBLUP largely revolves around the genetic and residual covariance matrices of the observed secondary data. We can obtain estimates of these matrices by using REML (restricted maximum likelihood) and the kinship matrix $\mathbf{K}$ for a small number of features. When the number of secondary features $p$ exceeds $5\text{-}10$, however, the computational complexity quickly becomes prohibitive and REML estimates can no longer be obtained (Zhou and Stephens, 2014). As this is usually the case for high-dimensional HTP data, we use data containing genotypic replicates and correct for experimental design factors and spatial trends to obtain pseudo-CRD (pseudo completely randomized design) data as a pre-processing step. Note again that in this case, assuming balanced data, the total number of records is simply the number of genotypes times the number of replicates, i.e., $n=n_{g}r$. We then assume that $\mathbf{Z}\mathbf{K}\mathbf{Z}^{\top}=\mathbf{Z}\mathbf{Z}^{\top}$, i.e., $\mathbf{K}=\mathbf{I}_{n_{g}}$, and obtain estimates for $\bm{\Sigma}_{ss}^{g}$ and $\bm{\Sigma}_{ss}^{\epsilon}$ using sums of squares as measures for between and within-genotype variation (Ott and Longnecker, 2015):
(4) $\begin{split}\mathbf{MS}_{ss}^{g}&=(\bar{\mathbf{Y}}_{s}^{g}-\bar{\mathbf{Y}}_{s})^{\top}(\bar{\mathbf{Y}}_{s}^{g}-\bar{\mathbf{Y}}_{s})(n_{g}-1)^{-1}\\ \mathbf{MS}_{ss}^{\epsilon}&=(\bar{\mathbf{Y}}_{s}^{g}-\mathbf{Y}_{s})^{\top}(\bar{\mathbf{Y}}_{s}^{g}-\mathbf{Y}_{s})(n-n_{g})^{-1}\\ \hat{\bm{\Sigma}}_{ss}^{g}&=(\mathbf{MS}_{ss}^{g}-\mathbf{MS}_{ss}^{\epsilon})r^{-1}\\ \hat{\bm{\Sigma}}_{ss}^{\epsilon}&=\mathbf{MS}_{ss}^{\epsilon}\text{,}\end{split}$
where $\mathbf{Y}_{s}$ is the $n\times p$ training set data matrix containing the secondary features. Overall means and genotypic means are located in $\bar{\mathbf{Y}}_{s}$ and $\bar{\mathbf{Y}}_{s}^{g}$, respectively. A well-known issue with the estimation of such covariance matrices using multivariate sums of squares is the fact that sufficient within-genotype variation can result in $\hat{\bm{\Sigma}}_{ss}^{g}$ having negative eigenvalues (Meyer and Kirkpatrick, 2010) as it is not a moment matrix but rather the difference of two moment matrices (as opposed to $\hat{\bm{\Sigma}}_{ss}^{\epsilon}$ which is a moment matrix and therefore at worst positive-semidefinite if $p>n$). We deal with this issue by simply finding the nearest positive-definite matrix using Higham’s algorithm (Higham, 2002). From here on, $\hat{\bm{\Sigma}}_{ss}^{g}$ refers to this nearest positive definite matrix.
We emphasize that the assumption of independent genetic effects is only made for the genetic and residual covariance estimators in equation (4). We use a marker-based kinship matrix in the final genomic prediction step. The assumption of diagonal kinship ($\mathbf{K}=\mathbf{I}_{n_{g}}$) in (4) simply means that the subsequent factor analysis is based on the covariance between the total (rather than only the additive) genetic effects.
2.2.3. Redundancy filtering (step II)
Given the fact that multicollinearity is common in HTP data, some of the $p$ secondary features will likely show near perfect genetic correlation and are therefore redundant. We filter the genetic correlation matrix using the algorithm described by Peeters et al. (2019) with an absolute genetic correlation threshold $\tau$ (i.e., after filtering, no two features will have an absolute genetic correlation $|\rho^{g}|$ equal to, or exceeding $\tau$). This filtering step produces $\hat{\mathbf{R}}_{ss}^{g*}\in\mathbb{R}^{p^{*}\times p^{*}}$ containing only the $p^{*}$ secondary features that pass filtering ($p^{*}\leq p$). We subset the residual correlation matrix and data matrix so they too only contain the remaining $p^{*}$ features.
2.2.4. Regularization (step III)
The filtered estimates $\hat{\mathbf{R}}_{ss}^{g*}$ and $\hat{\mathbf{R}}_{ss}^{\epsilon*}$ may still be ill-conditioned or singular which is why regularization is required. To achieve this, we use penalized maximum likelihood forms in the remainder of the glfBLUP method (Peeters et al., 2019; van Wieringen and Peeters, 2016):
where $\vartheta_{g},\vartheta_{\epsilon}\in(0,1]$. We find the optimal values for the penalty parameters $\vartheta_{g}^{\ddagger}$ and $\vartheta_{\epsilon}^{\ddagger}$ through $K$-fold cross-validation where we split the training data by dividing the $n_{g}$ genotypes over $K$ disjoint folds and minimize the following expression using Brent’s algorithm (Brent, 1973):
where $\left(\mathbf{R}_{ss}^{*}(\vartheta)\right)_{\neg k}$ is the penalized genetic or residual correlation matrix given the corresponding penalty $\vartheta$ estimated using the $K-1$ folds. The non-penalized correlation matrix $(\mathbf{R}_{ss}^{*})_{k}$ is estimated using the left-out fold. Each left-out fold is weighed by $w_{k}$, which is equal to the number of genotypes or individuals in that fold for the genetic and residual correlation matrix, respectively. See Peeters et al. (2019) for further details. We use an efficient implementation to minimize (6) which results in iterations of complexity $O(p^{*})$ as opposed to $O(p^{*3})$ due to the repeated inversion of $\mathbf{R}_{ss}^{*}(\vartheta)_{\neg k}$ required in the naive minimization. This leads to faster optimization, especially if optimal penalty values are close to the bounds. Details regarding our implementation can be found in SM Section 3. Note that both $\hat{\mathbf{R}}_{ss}^{g*}(\vartheta_{g}^{\ddagger})$ and $\hat{\mathbf{R}}_{ss}^{\epsilon*}(\vartheta_{\epsilon}^{\ddagger})$ balance the unbiased but high-variance estimated correlation matrices with the biased but stable identity matrix, optimizing the bias-variance trade-off. While our approach uses the identity matrix, we may use another target matrix $\mathbf{T}$ in (5) instead if prior information is available in the form of, for example, covariance estimates from earlier trials. The only requirement is that $\mathbf{T}$ is a positive-definite correlation matrix.
2.2.5. Latent dimension (step IV)
In principle, the model specified in (3) is valid for any latent dimension $m\leq\frac{1}{2}\left[2p+1-(8p+1)^{1/2}\right]$ (Ledermann, 1937). We must specify an estimate for $m$ prior to fitting the factor model and obtaining an estimate for the other parameters in $\bm{\Theta}$, however. To do so we use a lower eigenvalue bound based on the Marchenko-Pastur distribution (Marčenko and Pastur, 1967):
$\omega_{+}=\left(1+\sqrt{\dfrac{p^{*}}{n_{g}}}\right)^{2}\text{,}$
where $n_{g}$ is the number of genotypes used to estimate the genetic covariance matrix as described in Section 2.2.2. We then determine $\tilde{m}$:
2.2.6. Factor analytic modeling (step V)
With regularized estimates for the correlation matrices and latent dimension available, we now proceed with the factor analytic dimension reduction of the redundancy filtered secondary data. In the context of plant breeding, we are mainly interested in genetic correlations and will therefore fit the factor model using a reduced form of the secondary data covariance matrix. Recall that the full covariance of the BLUEs given in Section 2.2.1 was $\bm{\Sigma}_{ss}^{p*}=\bm{\Sigma}_{ss}^{g*}+r^{-1}\bm{\Sigma}_{ss}^{\epsilon*}$, with stars added to denote the redundancy filtering. The reduced covariance that we use to fit the factor model is thus $\bm{\Sigma}_{ss}^{p*}-r^{-1}\bm{\Sigma}_{ss}^{\epsilon*}=\bm{\Sigma}_{ss}^{g*}=\mathbf{\Lambda}\mathbf{\Lambda}^{\top}+\mathbf{\Psi}$. Given the fact that maximum-likelihood estimation of $\bm{\Theta}$ is scale-free (Peeters, 2012), we opt to estimate the loadings and uniquenesses using correlation rather than covariance matrices. We denote these correlation-scale parameters by $\bm{\Theta}_{R}=\{m,\mathbf{\Lambda}_{R},\mathbf{\Psi}_{R}\}$. Furthermore, we recognize that the sample genetic correlation matrix $\hat{\mathbf{R}}_{ss}^{g*}$ is a sufficient statistic for the population correlation matrix $\mathbf{R}_{ss}^{g*}$, meaning the maximum likelihood estimate $\hat{\bm{\Theta}}_{R}$, given a latent dimension $\tilde{m}$, can be found by minimizing the following discrepancy function (Jöreskog, 1967):
(7) $F\left[\mathbf{R}_{ss}^{g*}(\bm{\Theta}_{R});\hat{\mathbf{R}}_{ss}^{g*};\tilde{m}\right]=\ln\big{|}\mathbf{R}_{ss}^{g*}(\bm{\Theta}_{R})\big{|}+\operatorname{tr}\left[\hat{\mathbf{R}}_{ss}^{g*}\mathbf{R}_{ss}^{g*}(\bm{\Theta}_{R})^{-1}\right]-\ln\big{|}\hat{\mathbf{R}}_{ss}^{g*}\big{|}-p^{*}\text{,}$
where we constrain $\mathbf{\Lambda}_{R}^{\top}\mathbf{\Psi}_{R}^{-1}\mathbf{\Lambda}_{R}$ to be a diagonal matrix with ordered diagonal elements for identification. We replace $\hat{\mathbf{R}}_{ss}^{g*}$ by the regularized estimate $\hat{\mathbf{R}}_{ss}^{g*}(\vartheta_{g}^{\ddagger})$ as minimizing (7) requires a positive-definite and ideally well-conditioned sample genetic correlation matrix (Peeters et al., 2019). We employ a Varimax rotation once we have obtained the estimates for $\mathbf{\Lambda}_{R}$ and $\mathbf{\Psi}_{R}$ to facilitate interpretability (Kaiser, 1958). In essence, the varimax rotation attempts to bring the loadings matrix close to having simple structure, i.e., a structure in which variables load highly on a single factor and have near zero loadings on the other factors. To avoid notational clutter we simply use $\hat{\mathbf{\Lambda}}_{R}$ to refer to either the rotated or non-rotated estimated loadings. The described factor analytic model is implemented using an efficient matrix-free estimation method (Dai et al., 2020).
2.2.7. Estimation of factor scores (step VI)
We would now like to obtain estimates for the factor scores $\bm{\xi}_{j}$. These factor scores are essentially estimates of the locations of genotypes in latent space, based on the many noisy but observable secondary feature measurements. They can thus be viewed as the unobserved phenotypes of a small number of latent traits that generate the observed phenotypes for the many secondary features. We will use the joint distribution of the secondary feature BLUEs and latent factors (9) in a slightly modified Thomson regression approach (Thomson, 1939) to obtain $\hat{\bm{\xi}}_{j}$. Recalling that we estimated $\bm{\Theta}_{R}$ on the correlation scale, we first convert $\hat{\mathbf{\Lambda}}_{R}$ and $\hat{\mathbf{\Psi}}_{R}$ back to the covariance scale:
$\hat{\mathbf{\Lambda}}=\mathbf{D}\hat{\mathbf{\Lambda}}_{R}\text{ and }\hat{\mathbf{\Psi}}=\mathbf{D}\hat{\mathbf{\Psi}}_{R}\mathbf{D}\text{, where }\mathbf{D}=(\hat{\bm{\Sigma}}_{ss}^{g*}\circ\mathbf{I}_{p^{*}})^{1/2}\text{.}$
We can now obtain the matrix $\mathbf{P}$ mapping the redundancy filtered secondary features to their projections on the lower-dimensional latent space through $\mathbb{E}(\bm{\xi}_{j}|\bar{\mathbf{y}}_{s(j)}^{*})$. To obtain this expectation, we first use standard covariance algebra to obtain the joint distribution of the secondary feature BLUEs and latent factors as
@@ 2.2.7. Estimation of factor scores (step VI)
(9) $\begin{bmatrix}\bar{\mathbf{y}}_{s(j)}^{*\top}\\ \bm{\xi}_{j}\end{bmatrix}\sim\mathcal{N}_{p+m}\left[\mathbf{0},\bm{\Sigma}(\bm{\Theta})_{\xi}^{y}\right]\text{,}$
@@
Given that the joint distribution (9) is normal, we can obtain the conditional expectation as $\bm{\Sigma}(\bm{\Theta})_{\xi y}\left[\bm{\Sigma}(\bm{\Theta})_{yy}\right]^{-1}\bar{\mathbf{y}}_{s(j)}^{*}$ (Anderson, 2003). Using the Woodbury matrix identity (Woodbury, 1950), we rewrite this last expression to produce:
@@
$\begin{split}\tilde{\bm{\xi}}_{j}&=\Big{[}\mathbf{I}_{\tilde{m}}+\hat{\mathbf{\Lambda}}^{\top}(\hat{\mathbf{\Psi}}+r^{-1}\hat{\bm{\Sigma}}_{ss}^{\epsilon*})^{-1}\hat{\mathbf{\Lambda}}\Big{]}^{-1}\hat{\mathbf{\Lambda}}^{\top}(\hat{\mathbf{\Psi}}+r^{-1}\hat{\bm{\Sigma}}_{ss}^{\epsilon*})^{-1}\bar{\mathbf{y}}_{s(j)}^{*\top}\\ &=\mathbf{P}^{\top}\bar{\mathbf{y}}_{s(j)}^{*\top}\text{.}\end{split}$
@@
(10) $\begin{split}\tilde{\bm{\Xi}}&=\bar{\mathbf{Y}}_{s}^{*}\left[\hat{\mathbf{\Psi}}+r^{-1}\hat{\bm{\Sigma}}_{ss}^{\epsilon*}(\vartheta_{\epsilon}^{\ddagger})\right]^{-1}\hat{\mathbf{\Lambda}}\left\{\mathbf{I}_{\tilde{m}}+\hat{\mathbf{\Lambda}}^{\top}\left[\hat{\mathbf{\Psi}}+r^{-1}\hat{\bm{\Sigma}}_{ss}^{\epsilon*}(\vartheta_{\epsilon}^{\ddagger})\right]^{-1}\hat{\mathbf{\Lambda}}\right\}^{-1}\\ &=\bar{\mathbf{Y}}_{s}^{*}\mathbf{P}\text{.}\end{split}$
Through this projection we thus obtain factor scores that are essentially posterior estimates of latent phenotypes given the observed data. Instead of a full Bayesian approach we use empirical Bayes by using the maximum likelihood estimates of the parameters associated with the factor model in (10). Furthermore, note that $\text{var}(\bar{\mathbf{y}}_{s(j)}^{*})\rightarrow\bm{\Sigma}_{ss}^{g*}$ as $r\rightarrow\infty$, and as a result the projection $\mathbf{P}$ approaches the matrix that would be obtained using the ordinary Thomson regression approach. We thus account for the extra uncertainty due to a limited number of genotypic replicates.